OpenStudy (anonymous):

PLEASE HELP Factor the polynomial by grouping 4x^3+7x^2+3x

OpenStudy (anonymous):

4x^3+7x^2+3x =x(4x^2+7x+3) =x(4x^2+4x+3x+3) (4*3=12,4+3=7) =x{4x(x+1)+3(x+1)} =x(x+1)(4x+3)